OpenStudy (emogirlatemoooocow):

What is the value of r in the equation r + 9.2 = 17.5?

OpenStudy (anonymous):

@EmogirlAtEmoooocow ..make r the subject of the equation and send the constants to the other side. r=17.5-9.2

OpenStudy (igreen):

To find the value of r, get r by itself on one side. To do that subtract 9.2 from both sides: r + 9.2 - 9.2 = 17.5 - 9.2 r = 17.5 - 9.2 r = 8.3

OpenStudy (anonymous):

do the substraction and you will get the value of r

OpenStudy (igreen):

Check: r + 9.2 = 17.5 8.3 + 9.2 = 17.5 17.5 = 17.5 So 8.3 is the correct answer.